OpenStudy (anonymous):

Is 7^-3/2 not equivalent to 1/^2sqrt7^3 could have sworn the negative makes it 1 over exponent changed into radical form..unless I'm not converting it to radical form properly..

jimthompson5910 (jim_thompson5910):

You are correct, \[\Large 7^{-\frac{3}{2}} = \frac{1}{\sqrt[2]{7^3}}\] or \[\Large 7^{-\frac{3}{2}} = \frac{1}{\sqrt{7^3}}\]